# Moving Out: Physically-grounded Human-AI Collaboration

### v2

@@ C.1 Hyperparameters
Appendix D Reward Setting
D.1 Dense Reward Setting
@@ B. Event-based Rewards: Agent Holds an Item
Agent successfully holds an item Default reward for picking up +0.5+0.5+ 0.5
Exception: If another agent is already holding other middle or large item at this time 0.5-0.5- 0.5 (total for this hold event)
Exception: If the item picked up was already located within a goal region 0.5-0.5- 0.5 (total for this hold event)
@@ C. Event-based Rewards: Agent Unholds an Item
Agent successfully unholds an item Item is released inside a goal region +0.5+0.5+ 0.5
Item is released not inside a goal region 0.5-0.5- 0.5
Exception: If another agent needs help, holding a large or middle item outside the goal region, at the moment of unholding. +0.5+0.5+ 0.5 (additive)
@@ D. Time-based Reward (Step Cost)
Each timestep Agent exists in the environment 0.01-0.01- 0.01
D.2 Does MAPPO work in Moving Out with sparse reward setting?
@@
Appendix E Comparative analysis of the Behaviors of BC and RL agents
@@
Appendix F Details of Evaluation Metrics
We assess human-AI collaboration in Moving Out using the following metrics.
Task Completion Rate (TCR) The proportion of items successfully moved to the goal regions, weighted by size:
$TCR=\frac{\sum w_{i}\mathbb{I}(o_{i}\text{ delivered})}{\sum w_{i}}$
where $w_{i}=1$ (small) or 222 (middle/large). Range: [0,1].
Normalized Final Distance (NFD) The distance between objects and the closest goal regions, recognizing partial progress as valuable.
@@ Appendix F Details of Evaluation Metrics
where $d_{i}^{\text{initial}}$ and $d_{i}^{\text{final}}$ are the object’s initial and final distances to the target.
Waiting Time (WT) The total time an agent waits for help.
@@
where $\mathcal{W}$ is the set of time intervals when an agent holds a middle or large object but waits for help. $t_{\text{start}}^{t}$ and $t_{\text{end}}^{t}$ mark the waiting period.
Action Consistency (AC) This metric quantifies how well two agents align their forces when moving a middle or large object, indicating coordination efficiency.
@@
where $\vec{f}_{1}^{t},\vec{f}_{2}^{t}$ are the forces applied by the agents at time $t$, $\vec{d}_{t}$ is the unit vector connecting their positions, and $T$ is the total timesteps. The value ranges from 0 to 1, with 1 indicating perfect alignment.
Appendix G Implementation Details
G.1 Environment Details
G.1.1 Observation Encoding
@@
  • Self: Position and angle, with angles $\theta$ represented using $[\cos\theta,\sin\theta]$. A boolean value indicates whether the agent is holding an item (True/False).
@@ G.1.2 Action Encoding
  • The grasping action: 111 means grasp or release, 0 means no change.
G.2 Baseline Details
  • Diffusion Policy: We follow the original implementation by Chi et al. [2023] for the model architecture, which employs a 1D U-Net to generate action sequences. The observation, prediction, and executable horizons are set to 2, 8, and 4, respectively. Training is performed using the Adam optimizer with 1k epochs, 1024 batch size, and 0.001 learning rate. The diffusion steps are 36. The grasp action is encoded by one-hot encoding.
@@ G.3 BASS Details
In behavior augmentation, we add noise with a mean of 0 and a standard deviation of 0.0020.0020.002. In recombination sub-trajectories, since two points in a continuous space are almost never the same, we set a tolerance value. We discretize the environment into a $\times 48$ 48 grid. If the robot’s start and end points are in the same grid cell, we treat them as the same point.
@@
Many maps have walls, so we cannot use Euclidean distance. To improve efficiency, we discretize the environment into a $\times 48$ 48 grid. We use the BFS algorithm to compute the distance from the item to the Goal Region.
@@
Appendix H Data Collection: Training Data
For Task 1, which focuses on human behavior diversity, we recruited 36 participants, forming 18 groups of two. Before data collection, each group underwent a 10-minute practice session to familiarize itself with the environment. The remaining 50 minutes were dedicated to data collection. Each pair played each map three times, then switched agents and played three more times, resulting in six demonstrations per map. If a group completed all maps, they contributed a total of 12×6=7212$\times$ 6=7212 × 6 = 72 human demonstrations. However, not all groups completed the full set, with some collecting only 3 to 5 demonstrations per map. Additionally, we removed low-quality demonstrations where performance was significantly poor. In total, we collected 1,000 valid human demonstrations for this task.
For Task 2, which evaluates adaptation to physical constraints, we worked with four expert players who were highly familiar with the environment. Each map had randomized object properties, ensuring variation in shape, size, and mass. Each map was played 60 times, resulting in 60×12=72060$\times$ 12=72060 × 12 = 720 human demonstrations.
Our data collection and human study process was approved by the Institutional Review Board (IRB) at the University of Virginia (UVA), under protocol IRB-SBS #6932. Participants were compensated based on the amount of data they contributed, receiving between \$15 to \$20 per hour.
Appendix I Computing Resources
I.1 Training
I.1.1 Behavior Cloning
@@ I.1.2 MAPPO
I.2 Inference Speed of DP/BASS
@@
Appendix J Human Study: Playing with Models
J.1 Human Study Procedure
@@
Appendix K Results of Different Methods with BASS
@@
Table 6: The table presents all experimental results for Task 1
@@
Appendix L Map Analysis
L.1 Coordination

### v3

@@ E.1 Hyperparameters
Appendix F Reward Setting
F.1 Dense Reward Setting
@@ B. Event-based Rewards: Agent Holds an Item
Agent successfully holds an item Default reward for picking up +0.5+0.5
Exception: If another agent is already holding other middle or large item at this time 0.5-0.5 (total for this hold event)
Exception: If the item picked up was already located within a goal region 0.5-0.5 (total for this hold event)
@@ C. Event-based Rewards: Agent Unholds an Item
Agent successfully unholds an item Item is released inside a goal region +0.5+0.5
Item is released not inside a goal region 0.5-0.5
Exception: If another agent needs help, holding a large or middle item outside the goal region, at the moment of unholding. +0.5+0.5 (additive)
@@ D. Time-based Reward (Step Cost)
Each timestep Agent exists in the environment 0.01-0.01
F.2 Does MAPPO work in Moving Out with sparse reward setting?
@@
Appendix G Comparative analysis of the Behaviors of BC and RL agents
@@
Appendix H Additional Validation of Behavior Augmentation
Behavior mismatch in sub-trajectory recombination. Our recombination strategy is explicitly designed to avoid the type of inconsistency that was described. Specifically, we only perform sub-trajectory swapping when the fixed agent (e.g., agent A) has the same start and end poses across two trajectories. This ensures that agent A is pursuing the same local goal in both cases, regardless of the specific behavior of the partner.
For example, suppose in trajectory $\tau_{1}$, agent A performs action sequence $a_{1}$ while agent B performs $b_{1}$, and in $\tau_{2}$, A performs $a_{2}$ while B performs $b_{2}$. If $a_{1}$ and $a_{2}$ share the same start and end states, we can create two new combinations: $(a_{1},b_{2})$ and $(a_{2},b_{1})$. These are valid because both $b_{1}$ and $b_{2}$ were originally compatible with different variants of A’s strategy toward the same goal. As such, swapping B’s behavior does not interfere with A’s intent. This preserves behavioral diversity while ensuring trajectory-level coherence.
Additional validation. To ensure consistency, we identify sub-trajectories with matching start and end poses, so that the recombined agent behaviors maintain the same intention and goal. We have validated this approach with over 99% success rate in producing physically valid trajectories. In addition, we also confirm that the diversity of behavior increases after augmentation (Entropy improves from 0.88 to 0.95).
Appendix I Details of Evaluation Metrics
To assess human-AI collaboration in Moving Out, we design metrics that go beyond final task success to capture the quality of physical collaboration. While prior works such as Overcooked-AI mainly rely on task completion, this is insufficient in physically grounded settings, where interactions involve continuous control, object dynamics, and force alignment. We therefore complement Task Completion Rate (TCR) with three additional metrics—Normalized Final Distance (NFD), Waiting Time (WT), and Action Consistency (AC)—each targeting a different aspect of collaboration under physical constraints.
Task Completion Rate (TCR).
TCR measures the proportion of objects successfully delivered to goal regions, weighted by size:
$TCR=\frac{\sum w_{i}\mathbb{I}(o_{i}\text{ delivered})}{\sum w_{i}},$
where $w_{i}=1$ (small) or $2$ (middle/large). Range: [0,1]. TCR captures the final outcome of collaboration, but by itself cannot distinguish between failed attempts with meaningful progress and those with no progress.
Normalized Final Distance (NFD).
NFD quantifies partial progress by measuring the reduction in object-goal distance:
@@ Appendix I Details of Evaluation Metrics
where $d_{i}^{\text{initial}}$ and $d_{i}^{\text{final}}$ are the object’s initial and final distances to the target. This is critical in physical environments where objects may get stuck due to collisions or narrow passages. A case with high NFD but low TCR indicates that agents made progress but failed to overcome physical constraints.
Waiting Time (WT).
WT captures how agents coordinate when joint effort is required:
@@
where $\mathcal{W}$ is the set of intervals when an agent holds a middle/large object but must wait for help. High WT may reflect either poor recognition of the need for help or inefficiency in navigating physical obstacles. Thus, it measures awareness and responsiveness in physically grounded collaboration.
Action Consistency (AC).
As illustrated in Fig. 12, AC measures how well two agents align their applied forces during joint manipulation:
@@
where $\vec{f}_{1}^{t},\vec{f}_{2}^{t}$ are the forces applied at time $t$, $\vec{d}_{t}$ is the unit vector connecting agent positions, and $T$ is the number of timesteps. This metric captures coordination quality: agents are most effective when their forces are aligned, and receive low scores when their efforts cancel out (e.g., one pushing forward, the other pulling back).
Figure 12: Example of action consistency (AC) calculation. Effective collaborative work receives high scores, while opposing forces canceling each other lead to low scores.
Together, these four metrics provide a comprehensive evaluation of human-AI collaboration in physically grounded tasks: TCR reflects task success, NFD measures partial progress under constraints, WT captures coordination in joint effort, and AC quantifies the efficiency of physical interaction. This combination moves beyond symbolic settings and offers a richer view of how collaboration unfolds in continuous, low-level environments.
Appendix J Implementation Details
J.1 Environment Details
J.1.1 Observation Encoding
@@
  • Self: Position and angle, with angles $\theta$ represented using $\cos\theta,\sin\theta]$. A boolean value indicates whether the agent is holding an item (True/False).
@@
  • The grasping action: $1$ means grasp or release, 0 means no change.
J.2 Baseline Details
  • Diffusion Policy: We follow the original implementation by [17] for the model architecture, which employs a 1D U-Net to generate action sequences. The observation, prediction, and executable horizons are set to 2, 8, and 4, respectively. Training is performed using the Adam optimizer with 1k epochs, 1024 batch size, and 0.001 learning rate. The diffusion steps are 36. The grasp action is encoded by one-hot encoding.
@@
In behavior augmentation, we add noise with a mean of 0 and a standard deviation of $0.002$. In recombination sub-trajectories, since two points in a continuous space are almost never the same, we set a tolerance value. We discretize the environment into a $48\times 48$ grid. If the robot’s start and end points are in the same grid cell, we treat them as the same point.
@@
Many maps have walls, so we cannot use Euclidean distance. To improve efficiency, we discretize the environment into a $48\times 48$ grid. We use the BFS algorithm to compute the distance from the item to the Goal Region.
@@
Appendix K Study on Action Sampling Times
We conducted an ablation on the number and strategy of action samples used in BASS. In our implementation, each candidate action is generated by independently sampling from the policy and partner predictor up to four times, producing four simulation rollouts. This setting provides a good balance between performance and efficiency, supporting real-time human evaluation at 10Hz.
To compare alternatives, we tested three strategies: 1) Independent 4× sampling (current setting); 2) 2×2 combination (two samples from each agent, combined into four rollouts); 3) 4×4 combination (four samples from each, combined into sixteen rollouts).
Task 1 Task 2
[…]
$\times2$ combination 0.4522 0.5261 0.3485 0.8462 0.4158 0.5137 0.3101 0.8460
$\times4$ combination 0.5161 0.5847 0.3390 0.8727 0.4396 0.5638 0.3087 0.8559
Table 8: Results of sampling and combination strategies.
Results are summarized in Table 8. The 2×2 strategy, despite using the same number of simulations as the independent setting, consistently underperforms. Independent sampling has higher chance to capture critical joint transitions, e.g., resolving a stuck state. The 4×4 combination achieves the best accuracy, but requires 16 rollouts and increases inference time from 69 ms to 210 ms, which disrupts real-time human evaluation at 10Hz.
We therefore adopt the independent 4× sampling scheme in BASS, as it balances accuracy with the real-time feasibility required for human-in-the-loop collaboration.
Appendix L BASS Module Analysis
L.1 Next-State Prediction Accuracy
We evaluate the accuracy of the next-state prediction module by computing the L2 distance between the predicted state and the ground-truth state from the oracle simulator. Two baselines are included: a GRU-based predictor and a random state generator.
L2 Norm (↓) Task 1 Task 2
BASS 0.0010 0.0028
GRU 0.0196 0.0331
Random States 2.7576 3.2594
Table 9: Next-state prediction accuracy across tasks. Lower is better.
These results show that BASS more accurately captures physical dynamics compared to both the GRU baseline and random guessing, supporting the effectiveness of the learned dynamics model in simulation and action selection.
L.2 Modeling Diverse Human Behaviors
To handle diverse human behaviors, our approach models partner actions as a conditional distribution learned from demonstrations. The latent dynamics model captures this diversity by representing multiple likely behaviors under the same state, instead of committing to a single mode.
We use a Diffusion Policy, which effectively models multimodal action distributions by sampling from different noise inputs [53, 17]. This enables the model to generate different possible partner responses, providing probabilistic reasoning that aligns with human intuition.
L.3 Partner Action Prediction Accuracy
Since the action space is continuous, prediction accuracy is evaluated using the relative error between predicted and ground-truth actions. With a 10% error tolerance, the predictor achieves an accuracy of 71.45%; relaxing the tolerance to 20% increases accuracy to 90.24%. These results indicate that the partner action predictor provides sufficiently accurate estimates to support effective next-state prediction and action selection within our framework.
L.4 Effect of Random Partner Actions
To assess the importance of accurate partner action prediction, we conduct an ablation where the partner’s actions are randomly sampled during the simulation step. Since our action selection process considers four candidates, random guessing introduces uncertainty and degrades the ability to make correct selections between predicted future states.
TCR (↑) Action Selection Accuracy (↑)
Task 1 Task 2 Task 1 Task 2
BASS (Ours) 0.5027 0.4348 0.6250 0.4870
BASS (Random partner action) 0.3966 0.3650 0.2542 0.2484
BASS w/ Oracle Simulator 0.5421 0.5334 1.0000 1.0000
Table 10: Impact of replacing the learned partner model with random actions. Both task performance (TCR) and action selection accuracy drop significantly.
L.5 Comparison with Alternative Next-State Prediction Models
To evaluate the effectiveness of our proposed dynamics model, we compare it with two alternatives: a GRU-based predictor and a qVAE model for next-state prediction. As shown in Table 11, our model significantly outperforms both in prediction accuracy and final task performance (NFD). The GRU baseline performs close to random guessing, indicating difficulty in learning accurate transition dynamics. The qVAE model performs slightly better, but still struggles, likely due to the large continuous state space, where discretized latent codes are insufficient to represent fine-grained physical interactions. These results highlight the importance of our autoencoder-based latent dynamics model in capturing physical transitions effectively.
NFD (↑) Prediction Accuracy (↑)
Task 1 Task 2 Task 1 Task 2
BASS (Ours) 0.5733 0.5535 0.6250 0.4870
BASS (GRU) 0.5048 0.4965 0.2487 0.2598
BASS (qVAE) 0.5113 0.5032 0.3102 0.2722
BASS w/ Oracle Simulator 0.5875 0.6209 N/A N/A
Table 11: Comparison of next-state prediction models. Our dynamics model outperforms GRU and qVAE in both prediction accuracy and task performance.
Appendix M Evaluation Protocol for Task 1
Human-based evaluation. Since Task 1 is designed to evaluate adaptation to unseen human behaviors, our primary setup requires agents to play with new human participants. While this is the most direct evaluation of human-AI collaboration, it raises concerns about reproducibility because new participants are required for each run.
Cross-group reproducible protocol. To address this, we design a reproducible protocol inspired by cross-model evaluation. We randomly split the human demonstrations into two groups, each containing data from different participants. Two models are trained separately on each group and then evaluated by playing with each other. This setup emulates collaboration with unseen partner behavior while remaining fully reproducible.
Results. Table 12 compares the performance of Diffusion Policy (DP) and our method (BASS) when paired with models trained on the same group vs. a different group. We also report the percentage drop (or increase) in performance when moving from same-group to cross-group evaluation. Results show that models perform better when paired with a model trained on the same group (more aligned behavior), but BASS consistently outperforms DP when paired with unseen human behaviors, demonstrating stronger generalization.
Setting Method TCR (↑) NFD (↑) WT (↓) AC (↑)
Same-Group DP 0.3233 0.5367 0.3789 0.8163
DP/BASS 0.3503 0.5724 0.3598 0.8337
Cross-Group DP 0.2563 (-20.72%) 0.4589 (-14.50%) 0.4249 (+12.15%) 0.7854 (-3.78%)
DP/BASS 0.3010 (-14.07%) 0.5197 (-9.22%) 0.3899 (+8.37%) 0.8099 (-2.86%)
Table 12: Cross-group evaluation protocol for Task 1. Performance drops (%) are measured relative to same-group evaluation.
Appendix N Why choose Normalized Final Distance in BASS?
We chose Normalized Final Distance (NFD) because it directly reflects task progress in physically grounded collaboration. When two agents move an object together, actions that successfully reduce the distance between objects and the goal region indicate effective cooperation, even when navigating around obstacles like walls. Thus, maximizing NFD considers both physical feasibility and cooperation efficiency.
We also experimented with a multi-objective scoring using both NFD and Action Consistency (AC), to encourage not only progress but also force alignment. The trade-off is shown below:
Task 1 NFD ($\uparrow$) AC ($\uparrow$)
BASS with NFD 0.5733 0.8615
BASS with NFD+AC 0.5683 0.9127
Table 13: Comparison of using NFD vs. NFD+AC as objectives.
Although this combined objective improved AC, we found that NFD dropped slightly. We chose to prioritize NFD in the paper for its ability to capture physical task progress.
Appendix O Behavior Augmentation Details
Our augmentation strategy involves two techniques:
By aligning the start and end of agent $i$’s sub-trajectory, the generated trajectories maintain temporal consistency for agent $i$ while introducing a different sequence. This approach enriches the training set with new, valid trajectories where agent $i$’s behavior is fixed and the partner’s varies.
Appendix P Computing Resources
P.1 Training
P.1.1 Behavior Cloning
@@
P.2 Inference Speed of DP/BASS
@@
Appendix Q Data Collection: Training Data
For Task 1, which focuses on human behavior diversity, we recruited 36 participants, forming 18 groups of two. Before data collection, each group underwent a 10-minute practice session to familiarize itself with the environment. The remaining 50 minutes were dedicated to data collection. Each pair played each map three times, then switched agents and played three more times, resulting in six demonstrations per map. If a group completed all maps, they contributed a total of $12\times 6=72$ human demonstrations. However, not all groups completed the full set, with some collecting only 3 to 5 demonstrations per map. Additionally, we removed low-quality demonstrations where performance was significantly poor. In total, we collected 1,000 valid human demonstrations for this task.
For Task 2, which evaluates adaptation to physical constraints, we worked with four expert players who were highly familiar with the environment. Each map had randomized object properties, ensuring variation in shape, size, and mass. Each map was played 60 times, resulting in $60\times 12=720$ human demonstrations.
Our data collection and human study process was approved by an Institutional Review Board (IRB). Participants were compensated based on the amount of data they contributed, receiving between \$15 to \$20 per hour.
Appendix R Human Study: Playing with Models
R.1 Human Study Procedure
@@
Appendix S Full Results for Task 1 AI-AI Collaboration
This section reports the complete experimental results for Task 1 under AI-AI collaboration.
Evaluation Protocol Method TCR (↑) TCR StdErr NFD (↑) NFD StdErr WT (↓) WT StdErr AC (↑) AC StdErr Seen Behaviors MLP 0.2126 0.0072 0.2987 0.0048 0.4896 0.0021 0.8013 0.0093 GRU 0.2369 0.0183 0.3011 0.0142 0.4975 0.0202 0.8151 0.0173 MAPPO 0.1929 0.0038 0.3182 0.0045 0.5766 0.0068 0.8097 0.0071 DP 0.3233 0.0279 0.5367 0.0151 0.3789 0.0167 0.8163 0.0162 DP/BASS 0.3503 0.0293 0.5724 0.0232 0.3598 0.0182 0.8337 0.0146 Unseen Behaviors MLP 0.1433 0.0061 0.2413 0.0033 0.5647 0.0031 0.7729 0.0090 GRU 0.1638 0.0092 0.2453 0.0026 0.5758 0.0065 0.7830 0.0037 MAPPO 0.1635 0.0067 0.2808 0.0037 0.6379 0.0013 0.7858 0.0050 DP 0.2563 0.0152 0.4589 0.0177 0.4249 0.0136 0.7854 0.0041 DP/BASS 0.3010 0.0223 0.5197 0.0361 0.3899 0.0245 0.8099 0.0179 Play with Human DP 0.3855 0.0512 0.5547 0.0432 0.4886 0.0457 0.8054 0.0129 DP/BASS 0.6512 0.0717 0.7053 0.0459 0.3364 0.0481 0.9124 0.0113
Table 14: Task 1 results under seen and unseen human behaviors, and with real human partners.
Appendix T Results of Different Methods with BASS
@@
Table 15: The table presents all experimental results for Task 1 in seen behaviors.
@@
Appendix U Map Analysis
Our 12 maps are carefully designed to target specific collaboration modes, coordination, awareness, and action consistency, while ensuring that existing AI agents (e.g., MAPPO, Diffusion Policy) can perform some tasks but still exhibit clear limitations. This balance is essential: overly difficult maps with long paths or dense obstacles may yield near-zero performance for all agents, making it impossible to evaluate various aspects of human-AI collaboration. Our map definition is already based on a structural format, e.g., JSON, allowing easy modification, reuse of modules, and procedural generation for scalability.
U.1 Coordination